# Large-scale Long-tailed Disease Diagnosis on Radiology Images

### v2

@@ 6.2 Evaluation on Rad3D-DiagDS
Discussion on ROC Curves. The solid lines in the Figure 7 represents the median AUC value. This value is derived from a process where, for each category, 1000 random samples are taken and the median AUC value is calculated. This procedure is repeated multiple times (1000), with the final curve representing the median of these values. The ROC curve is illustrated by thess solid lines. Accompanying each solid lines is shaded area, which denotes the 95% confidence interval (CI). For each type of implementation, this shaded area is significantly higher than the portion above the solid line. This implies that the AUC value represented by the solid line is generally higher than the class-wise average across different data splits. As shown in the figure, the highest AUC value has reached 0.96, and the lowest value is 0.86, this observed pattern suggests, there exists a few categories that are more challenging to learn than others.
@@
    The numbers are borrowed from the referred papers. We are strictly align with them in train and test split.
6.3 External Evaluation
In this section, we explore the transferring ability of our final model, where we finetune the model on different downstream datasets across various imaging modalities, anatomies and target classes. As shown in Table 6.2, we can see significant performance improvement on various datasets with different data portions, compring to models trained from scratch. Additionally, in most cases, our model can also surpass former SOTAs significantly without any task-specific designs, e.g., architectures, loss functions, demonstrating that our RP3D-DiagDS can also serve as a superior large-scale supervised pre-training dataset for medical domain.
7 Clinical Impact
In this paper, we consider the problem of multi-modality, multi-anatomy, long-tailed disease diagnosis, which has great practical clinical meanings.
First, in clinical practice, patients may get multiple radiologic examinations during their long-term treatment progress from various medical departments. However, existing work on disease classification accepting one image scan, can hardly handle such circumstance, hindering the development of accurate and comprehensive diagnostic models. Our work compensate this and better meet the practical scenarios.
Second, comparing with existing diagnosis works that focus on common general disease classes, our efforts on long-tailed rare classes are more vital for clinical usage. For common diseases, AI diagnosis system, usually, can only help accelerate diagnosis procedure for clinicians, while the hint on rare classes is also critical.
Third, our final model can support finetuning on external diagnosis task. In clinical application, local medical centers can leverage this superiority even with a few training samples. This is remarkable especially considering that, for rare diseases, only a very few cases could be accessible in reality.
8 Limitation

### v3

@@ 2 RESULTS
Analysis on ROC Curves. The solid lines in the Figure 3 represents the median AUC value. This value is derived from a process where, for each category, 1000 random samples are taken and the median AUC value is calculated. This procedure is repeated multiple times (1000), with the final curve representing the median of these values. The ROC curve is illustrated by the solid lines. Accompanying each solid line is a shaded area, which denotes the 95% confidence interval (CI). For each type of implementation, this shaded area is significantly higher than the portion above the solid line. This implies that the AUC value represented by the solid line is generally higher than the class-wise average across different data splits. As shown in the figure, the highest AUC value has reached 0.97, and the lowest value is 0.91, this observed pattern suggests, that there exists a few categories that are more challenging to learn than others.
Analysis on Normal/Abnormal Diagnosis. In clinical scenarios, the proportion of normal cases is significantly higher compared to abnormal cases. Therefore, it is crucial to highlight the evaluation of the model’s ability to distinguish between normal and abnormal cases. As shown in Figure 5a, our model achieves an AUC of 82.61% for such diagnosis, given the significant disease variations in abnormal cases. Additionally, we also split out six test sub-sets based on the anatomy of cases, including Head and Neck, Chest, Abdomen and Pelvis, Upper Limb, Lower Limb, Spine, and Breast. In each sub-set, all cases are from the same anatomy, thus we can track the performance among a certain anatomy region. As shown in Figure 5a, the results on each anatomy are generally consistent, indicating that our model robustly demonstrates superior performance across different anatomical regions, without showing notable biases on any specific area.
Analysis on Prediction Probability Distribution. To get a deeper and more intuitive assessment of the model’s capabilities, we analyzed its performance across different disease classes. Given the large number of diseases, we pick a total of six diseases, as shown in Figure 4, where we demonstrate the prediction probability distribution on cases in the test set for a certain class. All cases are categorized by their related anatomies, with positive cases corresponding to the considered disease class and negative cases divided into two types: intra-negative cases from the same anatomy, and inter-negative cases from other anatomies. Horizontal dashed lines in the bar plot denotes the threshold, with positive and negative cases from the same anatomy colored in red and blue, and negative cases from other anatomies in different shades of gray. Ideally, positive cases should have high scores, while both types of negative cases should have low scores. Therefore, the difference between the actual distribution and the ideal situation can effectively illustrate the model’s performance on the disease. As shown by the results, for the success classes, the model can distinguish both the intra-negative and inter-negative cases against the positive, while for the ordinary classes, the model may be confused with some intra-negative and positive cases. In the failure class, the model cannot pick out the positive cases. These results demonstrate that our internal evaluation requires the model to perform a two-stage differential diagnosis. The first step is to locate the possible diseases through the shooting anatomy, etc., and then capture the detailed lesion patterns features to make further diagnosis.
Figure 4: The prediction probability distribution for different classes and saliency map visualization. a Based on the anatomies of each case, we split them into positive cases, intra-negative cases which are located in the same anatomy as the positive ones, and inter-negative cases which are located in other anatomies. The classification threshold score in the figure denotes the final comparison bar to transform the soft probabilities into binary true/false diagnosis results. The first three probability distribution figures depict the distributions of three relatively successful classes, where the model can clearly distinguish the inter-negative cases and the intra-negative cases are more confusing. We then show two ordinary classes. As shown by the distributions, most errors are caused by the intra-negative cases, and similarly, the inter-negative cases are easily dismissed as well. At last, we show a failure case where the model can hardly distinguish the positive and negative cases regardless of whether they are intra-negative or inter-negative. b Saliency map of the key frames. Red indicates the areas that the model focuses on when inferring the corresponding disease category. This indicates that RadDiag is capable of accurately identifying the locations of lesions or abnormal regions.
Analysis on Saliency Map Visualization. Additionally, inspired by the proposal of explainable AI (XAI) [40], we utilize Score-CAM [74], a visual explanation method based on class activation maps, to visualize the areas that contribute more to make a decision on the target disease class. As shown in Figure 4b, we sample several cases from different modalities and anatomies to compute the saliency map. For instance, as depicted in Figure 4b(1) for “Ulnar styloid impaction syndrome”, our model can pinpoint the location of the ulna during inference for this category, without being distracted by other anomalies, such as rings worn on the four finger. In another case in Figure 4b(2), for the category of “Spondylodiscitis”, our model shows its ability to focus on the intervertebral disk, with heightened attention particularly on the region of abnormality. It proves our model’s precision in identifying the relevant anatomical structures for the disease categories, even amidst other potential distractions or anomalies within the same image, thus illustrating its robustness and specificity in medical disease diagnosis.
Evaluation on External Benchmarks
In this part, we further assess the transferring ability of our diagnosis model trained on RP3D-DiagDS. We adopt two settings to demonstrate the transferring ability to External Benchmarks of our model, i.e., zero-shot, and fine-tuning. First, in the zero-shot setting, we evaluate the effectiveness of our model for directly deployed in real clinical scenarios. In detail, as our dataset encompasses a wide range of disease categories affecting the entire body, it can be tested on most external datasets. More details on how we perform zero-shot transferring for our method and other baselines can be found in Section 4.3.3. We have adopted F1, MCC, and ACC for comparison. Second, our model can also serve as a pre-trained model and be fine-tuned on each downstream task to improve the final performance. Specifically, for the dataset with single image input, we simply adopt the pre-trained visual encoder module, i.e., discarding the fusion module. For the dataset with multi-image input, we inherit the entire model. In both cases, the final classification layer will be trained from scratch. In addition to using all available external training data, we also consider using 1%, 10%, and 30% portion data for few-shot learning.
Figure 5: Analysis on normal/abnormal diagnosis. a The bar plot shows the AUC, AP, and Recall@0.05 of the Normal class on different body regions in the Internal test set of RP3D-DiagDS. b The radar figure shows the AUC and AP performance comparison between RadIN, BiomedCLIP, and Ours on external clinical datasets at five different anatomies.
Zero-shot Results on Normal/Abnormal Diagnosis. In zero-shot assessment, we evaluate the transferring ability of our final model on normal/abnormal diagnosis for external dataset, e.g., Brain-Tumor (405 normal, 906 abnormal), POCUS (256 normal, 302 abnormal), CT-Kidney (1037 normal, 1453 abnormal), MURA (1667 normal, 1530 abnormal), and VinDr-Spine (1070 normal, 1007 abnormal), totaling 9634 cases. These datasets are derived from medical centers in different regions, thus is a good assessment of the model’s performance in distinguishing normal from abnormal cases under zero-shot conditions. We directly adopt the “normal” classifier head trained on our RP3D-DiagDS to obtain the prediction probability score and compare it with other foundation models, RadIN [53] and BiomedCLIP [82], in terms of AUC and AP on the five zero-shot datasets. As shown in Figure 5b, our model demonstrates significantly better results on all the five considered datasets, i.e., Brain-Tumor, POCUS, CT-Kidney, and MURA, compared to RadIN and BiomedCLIP. These findings show that our model consistently demonstrates superior diagnostic performance on zero-shot normal/abnormal diagnosis, across different anatomies under cross-center real clinical practices, significantly outperforming existing SOTA models.
Zero-shot Results on Fine-grained Diagnosis. In addition to the normal/abnormal diagnosis, as shown in Table 3, we also conduct zero-shot evaluation on fine-grained diagnosis tasks introduced by each representative external datasets, spanning six anatomies, including the head and neck, chest, breast, abdomen and pelvis, and spine. These datasets cover five modalities, i.e., CT, MRI, X-ray, Ultrasound, and Mammography, allowing for a comprehensive evaluation of our model’s capabilities. As baselines, we have included RadIN [53] and BiomedCLIP [82] for comparison in the zero-shot setting. The results demonstrate that our model outperforms the other two models in overall performance, indicating strong transfer learning capabilities. Interestingly, our model’s zero-shot performance can sometimes be even comparable to the results obtained by the model trained from scratch on a subset of the corresponding dataset, demonstrating the superiority of our RadDiag for diagnostic capabilities, which cannot be achieved by former foundation models.
Table 3: Zero-shot results on 6 external datasets. For our model, we avoid exposing to these external data for zero-shot experiments and use the threshold corresponding to the highest F1 score on the evaluation set. For RadIN and BiomedCLIP, we use the threshold corresponding to the highest F1 score on each individual external dataset. We also report the supervised training results with partially down-stream task external data, denoting as “S.T.”, for demonstrating what diagnosis performance our model can achieve zero-shotly. We have evaluated F1, MCC and ACC, and all the results reported in the table are the averaged values across all classes. In the table, "US" represents ultrasound and "MG" represents mammography.
Dataset Anatomy Modality Dim Method External Data F1 $\uparrow$ MCC $\uparrow$ ACC $\uparrow$
Brain-Tumor Head&Neck MRI 2D RadIN Zero-shot 22.29 3.23 63.63
BiomedCLIP 56.25 32.81 57.21
Ours 61.09 47.37 73.38
S.T. n=571(10%) 61.77 46.69 72.17
POCUS Chest US Video RadIN Zero-shot 2.22 0.17 66.67
BiomedCLIP 50.33 11.73 45.82
Ours 65.02 47.97 79.84
S.T. n=223(1%) 64.20 48.31 80.58
CT-KIDNEY Abdom&Pelvis CT 2D RadIN Zero-shot 38.1 17.57 64.75
BiomedCLIP 49.38 29.58 72.21
Ours 56.50 43.18 69.96
S.T. n=995(10%) 55.73 42.28 66.95
MURA Limb X-ray 2D RadIN Zero-shot 31.26 4.37 52.94
BiomedCLIP 68.36 14.96 54.70
Ours 67.64 40.15 77.51
S.T. n=3680(10%) 66.71 41.98 78.32
VinDr-Spine Spine X-ray 2D RadIN Zero-shot 4.34 0.91 87.49
BiomedCLIP 26.49 12.79 61.3
Ours 25.93 19.40 83.07
S.T. n=84(1%) 24.73 20.61 84.46
TCGA Chest&Breast CT&MG 3D RadIN Zero-shot - - -
BiomedCLIP - - -
Ours 72.18 57.20 83.47
S.T. n=173(30%) 70.33 52.26 77.41
    All indicators are presented as percentages, with the % omitted in the table.
Finetuning Results. As shown in Table 4, while finetuning our model on external datasets, we see significant performance improvement on all 22 external datasets with different data portions, compared to models trained from scratch. In most cases, our model can even surpass the specialist SOTAs which are designed carefully for the targeting task, demonstrating that publicly shared medical data on the Internet is a tremendous and valuable resource that can serve as a superior large-scale supervised training dataset for the medical domain.
Table 4: The AUC score comparison on various external datasets. For each dataset, we experimented with different training data portions, denoted as 1% to 100% in the table. For example, 30% represents we use 30% of data in the downstream training set for finetuning our model or training from scratch. “SOTA” denotes the best performance of former works (pointed with corresponding reference) on the datasets. We mark the gap between ours and training from scratch on the subscript of $\uparrow$ in the table.
Dataset 1% 10% 30% 100% SOTA
Scratch Ours Scratch Ours Scratch Ours Scratch Ours
@@
[…]
    The numbers are borrowed from the referred papers.
  • The four SOTAs [54, 44, 81, 13] are relatively old. Thus, their network architecture is weaker than the classical 3D ResNet trained from scratch.
    All indicators are presented as percentages, with the % omitted in the table.
3 DISCUSSION
There have been a number of open-source datasets for disease diagnosis. Notably, large-scale datasets such as NIH ChestX-ray [75], MIMIC-CXR [39], and CheXpert [34] stand out for their comprehensive collection of annotated X-ray images, facilitating extensive research and advancements in automated disease classification. However, it’s important to note three key limitations. First, these open-source diagnosis datasets mainly consist of chest X-rays [59, 15, 75, 34, 51, 30, 37, 69, 26, 55, 39, 21], thus are 2D images. There are only a few 3D datasets available, with a limited number of volumes [3, 2, 4, 42, 52, 32, 12, 8, 11]. Second, a significant portion of these large-scale datasets focuses solely on binary classification of specific diseases [69, 26, 30, 37]. Third, existing large-scale datasets contain a broad range of disease categories in different granularities [15, 53]. For example, in PadChest [15], disorders exist at different levels, including infiltrates, interstitial patterns, and reticular interstitial patterns. Such variation in granularity poses a great challenge for representation learning. In summary, these datasets fall short of meeting real-world clinical needs, which often involve complex, multimodal, and multi-image data from a single patient. Therefore, constructing a dataset that mirrors the intricacies of actual clinical scenarios is necessary. In comparison, the large-scale radiology dataset we construct covers a large number of categories of diseases, including various common modalities.
The prevailing paradigm in earlier diagnostic models is a specialized model trained on a limited range of disease categories. These models focus on specific imaging modalities and are targeted towards particular anatomical regions. Specifically, ConvNets are widely used in medical image classification due to their outstanding performance, for example, [23, 71, 21] have demonstrated excellent results on identifying a wide range of diseases. Recently, Vision Transformer (ViT) has garnered immense interest in the medical imaging community, numerous innovative approaches  [22, 78, 61, 83, 77, 49, 28] have emerged, leveraging ViTs as a foundation for further advancements in this field.
The other stream of work is generalist medical foundation models  [73, 79, 84]. These models represent a paradigm shift in medical AI, aiming to create versatile, comprehensive AI systems capable of handling a wide range of tasks across different medical modalities, by leveraging large-scale, diverse medical data. MedPaLM M [73] reaches performance competitive with or exceeding the state-of-the-art (SOTA) on various medical benchmarks, demonstrating the potential of the generalist foundation model in disease diagnosis and beyond. RadFM [79] is the first medical foundation model capable of processing 3D multi-image inputs, demonstrating the versatility and adaptability of generalist models in processing complex imaging data. However, the development of GMAI models requires substantial computational power, which is often impractical for exploring sophisticated algorithms in academic labs.
In contrast, leveraging our proposed RP3D-DiagDS dataset, our final model RadDiag is capable of performing universal disease diagnosis and classification tasks within a constrained parameter scope, transcending the limitations of focusing on specific anatomies or disease categories. This capability offers a more versatile and efficient approach to healthcare management, resulting in significant clinical and research implications compared with precious AI models trained on former diagnostic datasets as discussed in Supplementary Section E. Our impacts and contributions can be summarized in the following three aspects:
Model Development for Case-Level Diagnosis. The RP3D-DiagDS dataset facilitates the creation of AI models capable of case-level diagnosis utilizing multi-modal and multi-scan inputs. This is a critical advancement since patients often undergo multiple radiological examinations across different medical departments during their treatment. Traditional disease classification models, which typically process single-image scans, are inadequate for such complex scenarios. Our dataset, with its case-level labels, promotes the development of comprehensive diagnostic AI systems, thereby addressing a significant gap in clinical practice.
Long-Tailed Rare Disease Diagnosis. Unlike existing datasets that primarily cover common diseases, the RP3D-DiagDS encompasses over 5,000 disorders, including a significant number of rare diseases. This focus on long-tailed rare disease diagnosis is particularly important for clinical applications, as AI systems can readily assist in diagnosing common diseases but often fall short in identifying rare conditions. Our dataset thus plays a crucial role in enhancing the diagnostic process for rare diseases, which are typically more challenging to diagnose due to their lower prevalence.
Value of Online Radiology Resources. The RP3D-DiagDS dataset, collected from the Internet, serves as a valuable resource for training disease diagnosis model, especially for local medical centers dealing with rare diseases. Given the scarcity of cases for such diseases, our dataset’s ability to support few-shot learning scenarios is of paramount importance. It demonstrates the value that, such publicly shared medical data on the Internet represents a colossal and high-quality resource, holding the potential to pave the way for constructing generalist models aimed at advancing AI4healthcare.